A 5.0-m-wide swimming pool is filled to the top. The bottom of the pool becomes completely shaded in the afternoon when the sun is 23 degrees above the horizon.

How deep is the pool? (in meters)
Answer #1

Start with Snell's Law

n1(sin angle 1) = n2 (sin angle 2)

1(sin 67) = 1.333(sin angle 2)

angle 2 = 43.67 degrees


Then apply the tangent function

tan (43.67) = 5/(the depth)

depth = 5.24 m